# Analog In-memory Training on General Non-ideal Resistive Elements: The Impact of Response Functions

### v3

@@ 3.2 Inexact Convergence of Analog SGD under generic devices
Assumption 3 requires that $W_{k}$ remains far from saturation points, which is a mild assumption in actual training. This paper considers the average gradient square norm as the convergence metric, given by $E_{K}:=\frac{1}{K}\sum_{k=0}^{K-1}\left\|\nabla f(W_{k})\right\|^{2}$. Now, we establish the convergence of Analog SGD.
@@
$E_{K}\leq$ $\ O\left(\sqrt{{\sigma^{2}}/{K}}+\sigma^{2}S^{\texttt{ASGD}}_{K}\right)$ (8)
@@
The proof of Theorem 2 is deferred to Appendix H. Theorem 2 suggests that the convergence metric $E_{K}$ is upper bounded by two terms: the first term vanishes at a rate of $O(\sqrt{{\sigma^{2}}/{K}})$, which matches the Digital SGD’s convergence rate [38] up to a constant; the second term contributes to the asymptotic error of Analog SGD, which does not vanish with the number of iterations $K$.
@@ 4 Mitigating Implicit Penalty by Residual Learning
The asymptotic error in Analog SGD is a fundamental issue that arises from the mismatch between the symmetric point and the critical point. An idealistic remedy for the inexact convergence is carefully shifting the weights to ensure the stationary point is close to a symmetric point. However, determining the appropriate adjustment is always challenging, as the critical point is difficult to pinpoint before the actual training. Therefore, an ideal solution to address this issue is to jointly construct a sequence with a predictable stationary point and a proper shift of the symmetric point.
Residual learning. Our solution overlaps the algorithmic stationary point and physical symmetric point on the special point 0. Besides the main analog array, $W_{k}$, we maintain another array, $P_{k}$, whose stationary point should be 0. A natural choice is the residual of the weight, $P^{*}(W)$, defined by the $P$ that minimizes the objective $f(W+\gamma P)$ with a non-zero $\gamma$. Additionally, the goal of the main array is to minimize the residual so that the model $W_{k}$ is approaching optimal. This process can be formulated as the following bilevel problem whose optimal points can be proved to be that of $f(W)$
$\textsf{Residual Learning}\quad\operatornamewithlimits{arg\,min}_{W\in{\mathbb{R}}^{D}}~\|P^{*}(W)\|^{2},\quad\operatornamewithlimits{s.t.}~P^{*}(W)\in\operatornamewithlimits{arg\,min}_{P\in{\mathbb{R}}^{D}}~f(W+\gamma P).$ (9)
Now we propose a gradient-based method to solve (9). The gradient of $f(W+\gamma P)$ with respect to $P$, given by $\nabla_{P}f(W+\gamma P)=\gamma\nabla f(W+\gamma P)$, is accessible with fair expense, enabling us to introduce a sequence $P_{k}$ to track the residual of $W_{k}$ by optimizing $f(W_{k}+\gamma P)$
@@
where $\bar{W}_{k}:=W_{k}+\gamma P_{k}$ is the mixed weight. We then derive the hyper-gradient of the upper-level objective. Notice $\nabla\|P^{*}(W)\|^{2}=2\nabla P^{*}(W)P^{*}(W)$. Assuming $W^{*}$ is the unique minimum of $f(\cdot)$, we know $P^{*}(W)$ satisfies $P^{*}(W)+W=W^{*}$. Taking gradient with respective to $W$ on both sides, we have $\nabla P^{*}(W)=-I$ and hence $\nabla\|P^{*}(W)\|^{2}=-2P^{*}(W)$. Approximating $P^{*}(W)$ by $P_{k}$, we reach the update of the main array.
@@
Featuring moving the residual $P_{k}$ to $W_{k}$, (11) is referred to as transfer process. The updates (10) and (11) are performed alternatively until convergence. It is noticeable that Residual Learning has the same recursion format as Tiki-Taka. However, as we will discuss in Section˜5, the proposed method can be naturally extended, and simulations in Section˜6 show that the extension of Residual Learning outperforms the extension of Tiki-Taka.
@@
If the optimal point $W^{*}$ exists and is unique, the solution of the lower-level objective has a closed form $P^{*}(W):=\frac{W^{*}-W}{\gamma}$. At that time, the upper-level objective equals $\|W^{*}-W\|^{2}$. However, the solutions of $f(\cdot)$ are non-unique in general, especially for non-convex objectives with multiple local minima. To ensure the existence and uniqueness of $W^{*}$, we assume the objective is strongly convex.
@@
Under the strongly convex assumption, the optimal point $W^{*}$ is unique. Although the requirement of strong convexity is non-essential by the development of bilevel optimization [40, 41, 42, 43], and the proof can be extended to more general cases, we leave it for future work.
@@ 5 Extension of Residual Learning: limited granularity and noisy IO
This section extends Residual Learning to more practical scenarios involving more hardware imperfections. To be specific, we consider the noisy IO and limited granularity as examples. We highlight that we are not trying to diminish the importance of imperfection, but we focus on two of the primary ones that are known to be crucial.
Noisy IO introduces noise during the reading of $P_{k+1}$ in the transfer process (11), given by $W_{k+1}=W_{k}+\beta(P_{k+1}+\varepsilon_{k})\odot F(W_{k})-\beta|P_{k+1}+\varepsilon_{k}|\odot G(W_{k})$ with a noise $\varepsilon_{k}$. It incurs the implicit penalty issues again, leading to a penalized upper-level objective $\|P^{*}(W)\|^{2}+\Sigma_{\varepsilon}\|W\|^{2}/2$, as claimed by Theorem˜1, where $\Sigma_{\varepsilon}$ is the standard variation of $\varepsilon_{k}$. To filter out the noise, we propose to use a digital buffer $H_{k}$ to take a moving average of noisy $P_{k+1}$ signals by
$H_{k+1}=$ $\ (1-\beta)H_{k}+\beta(P_{k+1}+\varepsilon_{k+1})$ (14)
@@
Furthermore, the optimization process suffers from a constant error in the order of $O(\Delta w_{\min})$ as the pulses are fired discretely, with each changing the weight by $O(\Delta w_{\min})$. To overcome these issues, we propose introducing a threshold mechanism to train in limited precision. Instead of transferring $H_{k+1}$ to $W_{k}$ at each iteration, the transfer is triggered only if an element in $H_{k+1}$ is greater than the response granularity $\Delta w_{\min}$. The proposed algorithms are referred to as Residual Learning v2.
@@ 6 Numerical Simulations
In this section, we verify the main theoretical results by simulations on both synthetic datasets and real datasets. We use the open source toolkit AIHWKIT (AIHWKIT[44] to simulate the behaviors of Analog SGD, Residual Learning (which reduces to Tiki-Taka). Each simulation is repeated three times, and the mean and standard deviation are reported. We consider two types of response functions in our simulations: power and exponential response functions with dynamic ranges $\tau,\tau]$, while the symmetric point is 0, as required by Corollary 1. More details, simulations, and ablation studies can be found in Appendix K.
MNIST FCN/CNN. We train a fully-connected network (FCN) and a convolutional neural network (CNN) on the MNIST dataset and compare the performance of Analog SGD and Tiki-Taka under various $\tau$ on power responses; see the results in Figure 4. By tracking residual, Tiki-Taka outperforms Analog SGD and reaches comparable accuracy with Digital SGD. For both architectures, the accuracy of Tiki-Taka drops by $1\%$. In contrast, Analog SGD takes a few epochs to achieve an observable accuracy increment in FCN training, rendering a slower convergence rate than Tiki-Taka. In CNN training, Analog SGD’s accuracy increases more slowly than Tiki-Taka does and finally gets stuck at about 80%. It is consistent with the theoretical claims.
CIFAR10/CIFAR100 ResNet. We fine-tune three ResNet models with different scales on CIFAR10/CIFAR100 datasets. The power response functions are used, whose results are shown in Table 1. The results show that the Tiki-Taka outperforms Analog SGD by about $1.0\%$ in most of the cases in ResNet34/50, and the gap even reaches about $7.0\%$ for ResNet18 training on the CIFAR100 dataset. On top of that, we also compare the proposed Residual Learning v2 and Tiki-Taka v2. Both of them outperform Residual Learning since they introduce a digital buffer to filter out the reading noise. However, Residual Learning v2 outperforms Tiki-Taka v2 on the CIFAR100 dataset, demonstrating the benefit from the bilevel formulation.
@@ 7 Conclusions and Limitations
This paper studies the impact of a generic class of asymmetric and non-linear response functions on gradient-based training in analog in-memory computing hardware. We first formulate the dynamic of Analog Update based on the pulse update rule. Based on it, we show that Analog SGD implicitly optimizes a penalized objective and hence can only converge inexactly. To overcome this issue, we propose a Residual Learning framework which solves a bilevel optimization problem. Explicitly aligning the algorithmic stationary point and physical symmetric point, Residual Learning provably converges to the optimal point exactly. Furthermore, we demonstrate how to extend Residual Learning to overcome the noisy reading and limited update granularity issues. The efficiency of the proposed method is verified through simulations. One limitation of this work is that the current analysis considers only the three hardware imperfections. While they are known to be crucial for analog training, it is also important to extend our convergence analysis and methods to more practical scenarios involving more imperfections in future work.
Appendix for “Analog In-memory Training on Non-ideal Resistive Elements:
Understanding the Impact of Response Functions”
@@ Appendix A Literature Review
Training on AIMC hardware. Analog training has shown promising early successes in certain tasks like face classification [25] and digit classification [26], where $1,000\times$ lower energy consumption compared with digital implementation is presented. Researchers are also exploring approaches to mitigate the impact of hardware non-idealities. For example, [27, 28] proposes to leverage the momentum technique to stabilize the training by reducing the noise. In order to overcome other potential non-idealities, a hybrid training paradigm is also being explored. [29] leverages the chip-in-the-loop technique to train models layer-by-layer, while [30] proposes to train the backbone in the digital domain and train the last layer in the analog domain. In general, these works have provided valuable insights into analog training, shedding light on many critical technical challenges. However, their focus has largely been on experimental and simulation aspects, with limited systematic and theoretical analysis of how specific imperfections affect the training process. In our paper, we present an alternative viewpoint and novel tools to explore the effects of non-idealities.
@@
However, the resistive element possessing symmetric updates may not be the best option in terms of manufacturing. For example, although ECRAM provides almost symmetric updates, it is still less competitive than ReRAM as ReRAM has a faster response speed and lower pulse voltage [49]. The suitability of the resistive elements is evaluated based on metrics across multiple dimensions, such as the number of conductance states, retention, material endurance, switching energy, response speed, manufacturing cost, and cell size. Among them, this paper is only interested in the impact of response functions in the training.
Imperfection of AIMC hardware. Besides the response functions, analog training suffers from all kinds of hardware imperfection, especially when the task’s scale increases, like asymmetric update [17, 19], reading/writing noise [18, 60, 61], device/cycle variations [62], non-linear current response due to IR drop [18, 6, 63]. This paper mainly focuses on asymmetric response functions. However, this paper is not trying to diminish the importance of other hardware imperfections but focuses on one of the primary ones that are known to be very important [19, 16].
Hardware-aware training. For inference on AIMC hardware purposes, models pretrained on digital hardware will be programmed on analog hardware. Due to the hardware imperfection, the pretrained models suffer from performance drops. Hardware-aware training (HWA) is a technique designed to bridge the gap between ideal pretrained models and non-ideal programmed models. In contrast to standard training methods, hardware-aware training explicitly incorporates device-specific imperfections, such as weight drift [20], device fail [64], bounded dynamic range [65], quantization error from ADC [66, 67, 68], device variation [69], and non-linear current output [70], into the training loop. By modeling these constraints during the training phase, the learned parameters become inherently more robust to real-world deployment conditions. It is worth highlighting that HWA is still performed on digital hardware, and the trained model will be programmed onto AIMC hardware. On the contrary, this paper considers a different and more challenging setting where the training is performed directly on analog hardware.
@@
Attributed to these efforts, analog training has empirically shown great promise in achieving a similar level of accuracy as digital training on chip prototype, with reduced energy consumption and training time [72, 73]. Simultaneously, the parallel acceleration solution with AIMC hardware is also under exploration [74]. Despite its good performance, it is still mysterious about when and why they work.
Theoretical foundation of gradient-based training. The closely related result comes from the convergence study of Tiki-Taka [21]. Similar to our work, they attempt to model the dynamic and provide the convergence properties of Analog SGD and Tiki-Taka. However, their work is limited to a special linear response function. Furthermore, their paper considers a simplified version of Tiki-Taka, with a hyper-parameter $\gamma=0$ (see Section˜4). As we will show empirically and theoretically, Tiki-Taka benefits from a non-zero $\gamma$. Consequently, We compare the results briefly in Table 2 and comprehensively in Appendix B.

### v4

@@ 3.2 Inexact Convergence of Analog SGD under generic devices
Assumption 3 requires that $W_{k}$ remains far from saturation points, which is a mild assumption in actual training. This paper considers the average squared norm of the gradient as the convergence metric, given by $E^{\texttt{ASGD}}_{K}:=\frac{1}{K}\sum_{k=0}^{K-1}\left\|\nabla f(W_{k})\right\|^{2}$. Now, we establish the convergence of Analog SGD.
@@
$E^{\texttt{ASGD}}_{K}\leq$ $\ O\left(\sqrt{{\sigma^{2}}/{K}}+\sigma^{2}S^{\texttt{ASGD}}_{K}\right)$ (8)
@@
The proof of Theorem 2 is deferred to Appendix H. Theorem 2 suggests that the convergence metric $E^{\texttt{ASGD}}_{K}$ is upper bounded by two terms: the first term vanishes at a rate of $O(\sqrt{{\sigma^{2}}/{K}})$, which matches the Digital SGD’s convergence rate [38] up to a constant; the second term contributes to the asymptotic error of Analog SGD, which does not vanish with the number of iterations $K$.
@@ 4 Mitigating Implicit Penalty by Residual Learning
The asymptotic error in Analog SGD is a fundamental issue that arises from the mismatch between the symmetric point and the critical point. An idealistic remedy for the inexact convergence is carefully shifting the weights to ensure the stationary point is close to a symmetric point. However, determining the appropriate shifting is challenging, as the critical point is unknown before training. Therefore, an ideal solution to address this issue is to jointly construct a sequence with a proper stationary point and a proper shift of the symmetric point.
Residual learning. Our solution overlaps the algorithmic stationary point and physical symmetric point on the special point 0. Besides the main analog array, $W_{k}$, we maintain another array, $P_{k}$, whose stationary point should be 0. A natural choice is the residual of the weight, $P^{*}(W)$, defined by the $P$ that minimizes the objective $f(W+\gamma P)$ with a non-zero $\gamma$. Notice that $P^{*}(W_{k})\to 0$ as $W_{k}\to W^{*}$. Additionally, the goal of the main array is to minimize the residual so that the model $W_{k}$ approaches optimality. This process can be formulated as the following bilevel problem, whose optimal points can be proved to be those of $f(W)$
$\textsf{Residual Learning}\quad\min_{W\in{\mathbb{R}}^{D}}~\|P^{*}(W)\|^{2},\quad\operatornamewithlimits{s.\!t.}~P^{*}(W)\in\operatornamewithlimits{arg\,min}_{P\in{\mathbb{R}}^{D}}~f(W+\gamma P).$ (9)
Now we propose a gradient-based method to solve (9). The stochastic gradient of $f(W+\gamma P)$ with respect to $P$, given by $\nabla_{P}f(W+\gamma P;\xi)=\gamma\nabla f(W+\gamma P;\xi)$, is accessible with fair expense, enabling us to introduce a sequence $P_{k}$ to track the residual of $W_{k}$ by optimizing $f(W_{k}+\gamma P)$
@@
where $\bar{W}_{k}:=W_{k}+\gamma P_{k}$ is the mixed weight. We then derive the hyper-gradient of the upper-level objective. Notice $\nabla\|P^{*}(W)\|^{2}=2\nabla P^{*}(W)P^{*}(W)$. Assuming $W^{*}$ is the unique minimum of $f(\,\cdot\,)$, we know $P^{*}(W)$ satisfies $\gamma P^{*}(W)+W=W^{*}$. Taking gradient with respective to $W$ on both sides, we have $\nabla P^{*}(W)=-\frac{1}{\gamma}I$ and hence $\nabla\|P^{*}(W)\|^{2}=-\frac{2}{\gamma}P^{*}(W)$. Approximating $P^{*}(W)$ by $P_{k}$ and absorbing $\frac{2}{\gamma}$ into the learning rate $\beta$, we reach the update of the main array
@@
Featuring moving the residual $P_{k}$ to $W_{k}$, (11) is referred to as transfer process. The updates (10) and (11) are performed alternatively until convergence. Tiki-Taka mentioned in [21] is the special case with linear response functions and $\gamma=0$.
@@
If the optimal point $W^{*}$ exists and is unique, the solution of the lower-level objective has a closed form $P^{*}(W):=\frac{W^{*}-W}{\gamma}$. At that time, the upper-level objective equals $\|W^{*}-W\|^{2}$. However, the solutions of $f(\,\cdot\,)$ are generally non-unique, especially for non-convex objectives with multiple local minima. To ensure the existence and uniqueness of $W^{*}$, we assume the objective is strongly convex.
@@
Under the strongly convex assumption, the optimal point $W^{*}$ is unique and hence the optimal solution of the lower-level problem in (9) is unique. Since the requirement of strong convexity is non-essential in the development of bilevel optimization [40, 41, 42, 43], we believe the proof can be extended to more general cases and will extend it for future work.
@@ 5 Extension of Residual Learning: limited granularity and noisy IO
This section extends Residual Learning to practical scenarios with additional hardware imperfections. To be specific, we consider the noisy IO and limited granularity as examples. We highlight that we are not trying to diminish the importance of imperfection, but rather focus on two of the primary ones known to be crucial.
IO of resistive crossbar arrays introduces noise during the reading of $P_{k+1}$ in the transfer process (11), given by $W_{k+1}=W_{k}+\beta(P_{k+1}+\varepsilon_{k})\odot F(W_{k})-\beta|P_{k+1}+\varepsilon_{k}|\odot G(W_{k})$ with a noise $\varepsilon_{k}$. It incurs the implicit penalty issues again, leading to a penalized upper-level objective $\|P^{*}(W)\|^{2}+\left\langle\Sigma_{\varepsilon},R_{c}(W)\right\rangle$, as claimed by Theorem˜1, where $\Sigma_{\varepsilon}={\mathbb{E}}[|\varepsilon_{k}|]$ is assumed to be a constant. To filter out the noise, we propose to use a digital buffer $H_{k}$ to take a moving average of noisy $P_{k+1}$ signals by
$H_{k+1}=$ $\ (1-\beta)H_{k}+\beta(P_{k+1}+\varepsilon_{k+1}).$ (14)
@@
Furthermore, the transfer process suffers from a constant error of $O(\Delta w_{\min})$ due to the discrete pulse firing, each of which changes the weight by $O(\Delta w_{\min})$. To overcome these issues, we propose introducing a threshold mechanism that does not transfer the entire $H_{k+1}$ to $W_{k}$ at each iteration, as in (15). Instead, we compute an intermediate value by $H_{k+\frac{1}{2}}=(1-\beta)H_{k}+\beta(P_{k+1}+\varepsilon_{k+1})$ first. At each coordinate $d$, if the value $|[H_{k+\frac{1}{2}}]_{d}|\geq\Delta w_{\min}$, one pulse will be fired to $W_{k}]_{d}$ and update the digital buffer by $H_{k+1}]_{d}=[H_{k+\frac{1}{2}}]_{d}-\Delta w_{\min}$ or $H_{k+1}]_{d}=[H_{k+\frac{1}{2}}]_{d}+\Delta w_{\min}$, where the sign of increment is determined by the sign of $H_{k+\frac{1}{2}}]_{d}$. Otherwise, no transfer is triggered if the intermediate value falls below the threshold, i.e., $H_{k+1}]_{d}=[H_{k+\frac{1}{2}}]_{d}$. The proposed algorithms are referred to as Residual Learning v2.
@@ 6 Numerical Simulations
In this section, we verify the main theoretical results by simulations on both synthetic datasets and real datasets. We use the open source toolkit AIHWKIT (AIHWKIT[44] to simulate the behaviors of Analog SGD, Residual Learning (which reduces to Tiki-Taka). Each simulation is repeated three times, and the mean and standard deviation are reported. We consider two types of response functions in our simulations: power and exponential response functions with dynamic ranges $\tau,\tau]$ and the symmetric point being 0, as required by Corollary 1. More details, simulations, and ablation studies can be found in Appendix K. The code of our simulations is available at github.com/Zhaoxian-Wu/analog-training.
FCN/CNN @ MNIST. We train a fully-connected network (FCN) and a convolutional neural network (CNN) on the MNIST dataset and compare the performance of Analog SGD and Tiki-Taka under various dynamic range $\tau$ on power responses; see the results in Figure 4. By tracking residual, Residual Learning outperforms Analog SGD and reaches comparable accuracy with Digital SGD. For both architectures, the accuracy of Residual Learning drops by $1\%$. In contrast, Analog SGD takes a few epochs to achieve a noticeable increase in accuracy in FCN training, rendering a slower convergence rate than Residual Learning. In CNN training, Analog SGD’s accuracy increases more slowly than Residual Learning, eventually settling at about 80%. It is consistent with the theoretical claims.
ResNet @ CIFAR10/CIFAR100. We fine-tune three ResNet models with different scales on CIFAR10/CIFAR100 datasets. The power response functions are used, whose results are shown in Table 1. The results show that the Tiki-Taka outperforms Analog SGD by about $1.0\%$ in most of the cases in ResNet34/50, and the gap even reaches about $7.0\%$ for ResNet18 training on the CIFAR100 dataset. On top of that, we also compare the proposed Residual Learning v2 and Tiki-Taka v2. Both of them outperform Residual Learning since they introduce a digital buffer to filter out the reading noise. However, Residual Learning v2 outperforms Tiki-Taka v2 on the CIFAR100 dataset, demonstrating the benefit from the bilevel formulation. Ablation study on $\gamma$. We conduct simulations to study the impact of mixing coefficient $\gamma$ in (10) on the CIFAR10 or CIFAR100 dataset in the ResNet training tasks. The results are presented in Figure 5, which shows that Residual Learning achieves a great accuracy gain from increasing $\gamma$ from 0 to 0.1, while the gain saturates from 0.1 to 0.4. Therefore, we conclude that Residual Learning benefits from a non-zero $\gamma$, and the performance is robust to the $\gamma$ selection.
Figure 5: The test accuracy of ResNet family models after 100 epochs trained by Residual Learning under different $\gamma$ in (10); (Left) CIFAR10. (Right) CIFAR100.
@@ 7 Conclusions and Limitations
This paper studies the impact of a generic class of asymmetric and non-linear response functions on gradient-based training in analog in-memory computing hardware. We first formulate the dynamics of Analog Update based on the pulse update rule. Based on it, we show that Analog SGD implicitly optimizes a penalized objective and hence can only converge inexactly. To overcome this issue, we propose a Residual Learning framework which solves a bilevel optimization problem. Explicitly aligning the algorithmic stationary point and physical symmetric point, Residual Learning provably converges to the optimal point exactly. Furthermore, we demonstrate how to extend Residual Learning to overcome the noisy reading and limited update granularity issues. The efficiency of the proposed method is verified through simulations. One limitation of this work is that the current analysis considers only the three hardware imperfections. While they are known to be crucial for analog training, it is also important to extend our convergence analysis and methods to more practical scenarios involving more imperfections in future work.
Appendix for “Analog In-memory Training on Non-ideal Resistive Elements: Understanding the Impact of Response Functions”
@@ Appendix A Literature Review
Training on AIMC hardware. Analog training has shown promising early successes in tasks such as face classification [25] and digit classification [26], achieving $1,000\times$ lower energy consumption than digital implementations. Researchers are also exploring approaches to mitigate the impact of hardware non-idealities. For example, [27, 28] proposes leveraging the momentum technique to stabilize training by reducing noise. To address other potential non-idealities, a hybrid training paradigm is also being explored. [29] leverages the chip-in-the-loop technique to train models layer-by-layer, while [30] proposes to train the backbone in the digital domain and train the last layer in the analog domain. In general, these works have provided valuable insights into analog training, shedding light on many critical technical challenges. However, their focus has largely been on experimental and simulation aspects, with limited systematic and theoretical analysis of how specific imperfections affect the training process. In our paper, we present an alternative viewpoint and novel tools to explore the effects of non-idealities.
@@
However, a resistive element with symmetric updates may not be the best option for manufacturing. For example, although ECRAM provides almost symmetric updates, it remains less competitive than ReRAM, which offers faster response speed and lower pulse voltage [49]. The suitability of the resistive elements is evaluated using metrics across multiple dimensions, including the number of conductance states, retention, material endurance, switching energy, response speed, manufacturing cost, and cell size. Among them, this paper is only interested in the impact of response functions in the training.
Imperfection of AIMC hardware. Besides the response functions, analog training suffers from all kinds of hardware imperfection, especially when the task’s scale increases, like asymmetric update [17, 19], reading/writing noise [18, 60, 61], device/cycle variations [62], non-linear current response due to IR drop [18, 6, 63]. This paper mainly focuses on asymmetric response functions. However, this paper is not trying to diminish the importance of other hardware imperfections but rather focuses on one of the primary ones known to be very important [19, 16].
Hardware-aware training. For inference on AIMC hardware purposes, models pretrained on digital hardware will be programmed on analog hardware. Due to hardware imperfections, the pretrained models suffer performance drops. Hardware-aware training (HWA) is a technique designed to bridge the gap between ideal pretrained models and non-ideal programmed models. In contrast to standard training methods, hardware-aware training explicitly incorporates device-specific imperfections, such as weight drift [20], device fail [64], bounded dynamic range [65], quantization error from ADC [66, 67, 68], device variation [69], and non-linear current output [70], into the training loop. By modeling these constraints during training, the learned parameters become inherently more robust to real-world deployment conditions. It is worth highlighting that HWA is still performed on digital hardware, and the trained model will be programmed onto AIMC hardware. On the contrary, this paper considers a different, more challenging setting in which training is performed directly on analog hardware.
@@
Attributed to these efforts, analog training has empirically shown great promise, achieving accuracy comparable to that of digital training on chip prototypes while reducing energy consumption and training time [72, 73]. Simultaneously, the parallel acceleration solution with AIMC hardware is under exploration [74]. Despite its good performance, it remains mysterious when and why the analog training works.
Theoretical foundation of gradient-based training. The closely related result comes from the convergence study of Tiki-Taka [21]. Similar to our work, they attempt to model the dynamics and provide the convergence properties of Analog SGD and Tiki-Taka. However, their work is limited to a special linear response function. Furthermore, their paper considers a simplified version of Tiki-Taka, with a hyper-parameter $\gamma=0$ (see Section˜4). As we will show empirically and theoretically, Tiki-Taka benefits from a non-zero $\gamma$. Consequently, We compare the results briefly in Table 2 and comprehensively in Appendix B.